Find f + g and f – g.

f(x)=3x squared plus 2x, g(x)=-x suared plus 4x

Answer by edjones(8007)

f(x)=3x^2+2x
g(x)=-x^2+4x
f+g=2x^2+6x
f-g=4x^2-2x
.
Ed